# Beyond KAN: Introducing KarSein for Adaptive High-Order Feature Interaction Modeling in CTR Prediction

### v3

@@ Abstract
Modeling feature interactions is crucial for click-through rate (CTR) prediction, particularly when it comes to high-order explicit interactions. Traditional methods struggle with this task because they often pre-define a maximum interaction order, which relies heavily on prior knowledge and can limit the model’s effectiveness. Additionally, modeling high-order interactions typically leads to increased computational costs. Therefore, the challenge lies in adaptively modeling high-order feature interactions while maintaining efficiency. To address this issue, we introduce Kolmogorov-Arnold Represented Sparse Efficient Interaction Network (KarSein), designed to optimize both predictive accuracy and computational efficiency. We firstly identify limitations of directly applying Kolmogorov-Arnold Networks (KAN) to CTR, and then introduce KarSein to overcome these issues. It features a novel architecture that reduces the computational costs of KAN and supports embedding vectors as feature inputs. Additionally, KarSein employs guided symbolic regression to address the challenge of KAN in spontaneously learning multiplicative relationships. Extensive experiments demonstrate KarSein’s superior performance, achieving significant predictive accuracy with minimal computational overhead. Furthermore, KarSein maintains strong global explainality while enabling the removal of redundant features, resulting in a sparse network structure. These advantages also position KarSein as a promising method for efficient inference  [^1] https://github.com/Ancientshi/KarSein.

### v4

# CTR-KAN: KAN for Adaptive High-Order Feature Interaction Modeling

@@ Abstract
Modeling high-order feature interactions is critical for click-through rate (CTR) prediction, yet traditional approaches often face challenges in balancing predictive accuracy and computational efficiency. These methods typically rely on pre-defined interaction orders, which limit flexibility and require extensive prior knowledge. Moreover, explicitly modeling high-order interactions can lead to significant computational overhead. To tackle these challenges, we propose CTR-KAN, an adaptive framework for efficient high-order feature interaction modeling. CTR-KAN builds upon the Kolmogorov-Arnold Network (KAN) paradigm, addressing its limitations in CTR prediction tasks. Specifically, we introduce key enhancements, including a lightweight architecture that reduces the computational complexity of KAN and supports embedding-based feature representations. Additionally, CTR-KAN integrates guided symbolic regression to effectively capture multiplicative relationships, a known challenge in standard KAN implementations. Extensive experiments demonstrate that CTR-KAN achieves state-of-the-art predictive accuracy with significantly lower computational costs. Its sparse network structure also facilitates feature pruning and enhances global interpretability, making CTR-KAN a powerful tool for efficient inference in real-world CTR prediction scenarios. The source code is available at  [^1] https://github.com/Ancientshi/KarSein.